Q: A consumer's utility function is given by the expression: = (0.6X0.5+ 0.4Y0.5)2.

  • Determine the marginal utility functions for each commodity. Does marginal utility decrease when consumption increases?
  • Assuming that the price of good is Rs 15 and the price of Y is Rs 6, write the equation of the budget line and plot it when income is Rs 450. What is its slope? What does it indicate?
  • Calculate the marginal rate of substitution of Y for X and interpret its economic meaning. Write the equation showing the consumer's equilibrium condition.
  • Obtain the equilibrium values of X and Y.
  • Find the expressions for change in MUx due to an increase in Y and change in MUy due to an increase in X.
1
Expert's answer
2021-03-02T07:52:45-0500

u=(0.6X0.5+0.4Y0.5)2


a. MUx=2(0.6X0.5+0.4Y0.5)(0.3X-0.5)

MUy=2(0.6X0.5+0.4Y0.5)(0.2Y-0.5)

Marginal utility decreases as consumption of the commodity increases.


b. if price is of X is RS15 ; price of Y is RS6 ; and income is RS450

then the equation of the of the consumer is given as:

"450=15X+6Y"

and the equation of the budget line is

"Y=75-2.5X" 

and the slope = -2.5

The slope shows the ratio of the prices


c. MRSy for x "\frac{2(0.6X^{0.5}+0.4Y^{0.5})(0.3X^{-0.5})}{2(0.6X^{0.5}+0.4Y^{0.5})(0.2Y^{-0.5})}"

"\\frac{0.3X^{-0.5}}{0.2Y^{-0.5}}\n\n\u200b"

"\frac{3}{2}{(\frac{Y}{X})}^{0.5}"

The consumer's equilibrium condition is given as

"\frac{MUx}{MUy}=\frac{Px}{Py}"

"\frac{3}{2}{(\frac{Y}{X})}^{0.5} = \frac{15}{6}"


d. "\frac{3}{2}{(\frac{Y}{X})}^{0.5} = \frac{15}{6}"

"(\frac{Y}{X})^{0.5}=\frac{15}{6}×\frac{2}{3}"

"(\frac{Y}{X})^{0.5}=\frac{5}{3}"

"\frac{Y}{X}=(\frac{5}{3})^{2}"

Y = 2.78X

Substituting the value of Y in the budget line:

"450=15X+6Y"

"450=15X+6(2.78X)"

450=15X + 16.7X

450=31.7X

X= 14.2


Y= 2.78(14.2)

Y= 39.8

The equilibrium values for X and Y are (14.2,39.8)


e. i) "\Delta{MUx} =\frac{\delta{MUx}}{\delta{y}}{\Delta{Y}}"

"0.12({XY})^{-0.5}\Delta{Y}"


ii) "∆MUy=\frac{\delta{MUy}}{\delta{x}}\Delta{X}"

= 0.12"(XY)^{-0.5}\Delta{X}"